Suppose that the diameters of golf balls manufactured by a certain company are normally

distributed with mean of 1.96 inches and standard deviation of 0.04 in. A golf ball will be

considered defective if its diameter is less than 1.90 in. or greater than 2.04 in. What is the

percentage of defective balls manufactured by the company?


1
Expert's answer
2021-11-22T19:05:44-0500

Let x be the diameter of a random ball, than X ~ "N(1.96,0.04^2)" = 1.96 + 0.04N(0, 1)

Lets find the probability of the defective ball as 1 minus probability of a normal

"P(normal)=P(1.9<X<2.04)=P(1.9<1.96+0.04N(0, 1)<2.04)=P(-1.5<N(0, 1)<2)=P(N(0,1)<2)-P(N(0,1)<-1.5)=0.97725-0.06681=0.27333=0.91044\implies P(defective)=1-0.91044=0.08956"

So, the percentage of the defective balls is 100% * 0.08956 = 8.96%